OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!!! HELP!!!!!! What is 7x / 3.49y - 2z when x = 3.6 y = 100 and z = 1,200? *NO ANSWER CHOICES PROVIDED!*

OpenStudy (juscallmesteve):

-1677.94

OpenStudy (anonymous):

\[\frac{x }{3.49y} - 2z\] like that?

OpenStudy (anonymous):

Hi yeah Satellite, thanks! You too Steve!

OpenStudy (anonymous):

\[\frac{x }{3.49y} - 2z\\ \frac{3.6 }{3.49\times 100} - 2\times 1200\] and a calculator

OpenStudy (anonymous):

Thank you Satellite!

OpenStudy (juscallmesteve):

my math was wrong, 25.2/349 = 0.0722063 0.0722063-2400 = -2399.93

OpenStudy (anonymous):

yw

OpenStudy (juscallmesteve):

(7 x/3.49 y) - 2 z and got -1677.94 not sure why, maybe i set my variables incorrectly